a compound gear train in which, wheel A is fixed to the driving shaft, B-C are


compounded , D is idle, and E is fixed on the driven shaft. If gear A rotates at 1000 r.p.m find


the speed and direction of E. driven shaft.

1
Expert's answer
2021-11-11T05:17:02-0500

Drive shaft rotates at 1000rpm=Gear A

x+y=1000 this will be equation (i)

Gear E is fixed , "y+x*(\frac{T_AT_D}{T_CT_B})=0"

Taking TA=30, TB=32, TC=30, TD=22

"y+x*(\frac{30*22}{20*32})=0"

"y+x(\frac{66}{64})=0\implies y=-x(\frac{66}{64})"

Use in equation (i)

"x+(-x(\frac{66}{64}))=1000"

"\frac{64x-66x}{64}=1000"

-2x=64000

x=-32000

The required speed, y"=-x(\frac{66}{64})"

"=-(-32000)(\frac{66}{64})"

=33000rpm in the clockwise direction since it's positive.